# Integrating Dynamical Systems Modeling with Spatiotemporal scRNA-seq Data Analysis

### v1

@@ 3 Dynamic Modeling of Single-cell Transcriptomics
In this section, we describe the dynamical systems models for scRNA-seq datasets. We begin with methods for snapshot data, i.e. cells sequenced at a single time point, including pseudotime methods, discrete Markov Chain methods, and continuous RNA velocity methods and their extensions. Next, we review methods targeted for temporally resolved scRNA-seq data, majorly based on various formulations and extensions of optimal transport (OT) based methods.
@@ 3.1 Snapshot Single-cell RNA-seq
When analyzing such a snapshot of ”cell state ensembles”, several approaches have been developed to uncover the underlying dynamical processes. One popular type of method, pseudotime, ranks individual cells temporally based on the structure of the data manifold or prior biological knowledge. Other techniques focus on modeling stochastic dynamics over the point clouds of observed cells, yielding discrete random-walk analyses. Additionally, continuous differential equation models have been proposed to infer the data-generating process of snapshot scRNA-seq dataset, to use these models for future dynamical predictions. In the following sections, we will explore these methods in more detail.
@@ 3.1.1 Pseudotime Methods
Given snapshot data from single-cell sequencing, where the data matrix $\textbf{X}\in\mathbb{R}^{n\times d}$, pseudotime assigns a positive real number for each cell to reflect its order during a dynamical process. Let $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represent the state vector (such as mRNA expression) for the $i$-th cell. The pseudotime $t_{i}\in\mathbb{R}$ is then a mapping from the state vector $\mathbf{x}_{i}$ to a real number, i.e., $\mathbf{x}_{i}\mapsto t_{i}$.
Pseudotime can be viewed from two perspectives. First, the geodesic perspective considers that the cell’s state is constrained by a limited number of biological pathways, limiting the evolution of the state to a low-dimensional manifold embedded in the high-dimensional gene expression space. Given an initial state $\mathbf{x}_{0}$, the task of finding the mapping $\mathbf{x}_{i}\mapsto t_{i}$ becomes equivalent to determining the length of the evolutionary path between $\mathbf{x}_{0}$ and $\mathbf{x}_{i}$ along this manifold. Several tools, such as Monocle, Slingshot, DPT, and PAGA, have been developed based on this idea. Second, the entropy perspective recognizes that during natural biological development, as cells differentiate, they become more specialized and lose their potential for further differentiation.
Geodesic-based Pseudotime
Geodesic-based Pseudotime aims to reconstruct these trajectories by leveraging graph-based methods and principal curve algorithms (Deconinck et al., 2021). Two widely used approaches include Monocle (Qiu et al., 2017; Cao et al., 2019) and Slingshot (Street et al., 2018), along with numerous other methods. Monocle focuses on ordering cells using a minimum spanning tree and a refined PQ tree approach, and Slingshot can handle multiple lineages and smooths pseudotime across branching events. Both methods offer insights into cellular dynamics, helping to uncover the paths cells take through different states.
Monocle estimates pseudotime through two main steps: (1) ordering cells and (2) assigning pseudotime values. First, each cell’s state is reduced to a $d$-dimensional vector $\mathbf{x}_{i}\in\mathbb{R}^{d}$ using Independent Component Analysis (ICA). A complete graph is created where vertices represent cells, and edges are weighted by Euclidean distance. Cell ordering relies on the minimum spanning tree (MST) of this graph, which is refined using a PQ tree to mitigate noise from sequencing. The tree is constructed by first identifying the longest path (diameter path), classifying vertices as decisive or indecisive, and recursively building the tree by ordering decisive vertices and handling indecisive vertices through new P nodes. Once ordered, pseudotime is calculated as
$t(\mathbf{x}_{i})=t(\mathbf{x}_{\text{Parent}(i)})+\|\mathbf{x}_{i}-\mathbf{x}_{\text{Parent}(i)}\|,$
@@
Slingshot estimates pseudotime across multiple lineages. It begins by clustering cells into $K$ clusters and identifies lineages using the MST. After constructing the MST, a new lineage is formed at each branching point. Pseudotime is assigned using the principal curve algorithm, which involves projecting cells onto the curve, computing arc lengths, and smoothing iteratively. To handle inconsistent pseudotime across multiple lineages, Slingshot modifies the standard principal curve approach. It initializes the curve for each lineage through the centroids of its clusters, assigns weights to cells in multiple lineages based on projection distances, and constructs an average curve for smooth transitions at shared cell regions. The average curve is defined as
$\mathbf{c}_{\text{avg}}(t)=\frac{1}{M}\sum_{m=1}^{M}\mathbf{c}_{m}(t),$
where $\mathbf{c}_{m}$ is the principal curve of the $m$-th lineage at the branching point. The shrinkage process is defined as
$\mathbf{c}_{m}^{\text{new}}(t)=w_{m}(t)\mathbf{c}_{\text{avg}}(t)+(1-w_{m}(t))\mathbf{c}_{m}(t),$
@@
Entropy-based Pseudotime
One major challenge of the geodesic-based pseudotime is the appropriate determination of root cells, which often relies on prior biological knowledge. From a physical understanding, a cell’s pseudotime reflects the directionality of the underlying dynamical process, which the concept of entropy could quantify. Heuristically, higher entropy values typically indicate a more undifferentiated or pluripotent state where genes are more randomly expressed and the association between genes could be more prevalent. In comparison, lower entropy values suggest differentiated states, where the gene expression profile could be more concentrated on only a small number of pathways, and the gene interaction network could be more modular (Gandrillon et al., 2021, 2023). A proper entropy score based on such intuitions can thus be leveraged to estimate a cell’s relative position along developmental trajectories. Several methods have been developed from this perspective (Grün et al., 2016; Guo et al., 2017; Teschendorff and Enver, 2017; Shi et al., 2020; Jin et al., 2018; Liu et al., 2020).
As a simple implementation, the entropy for one cell $i$ is defined as (Grün et al., 2016)
@@
where $p_{ij}=\frac{X_{ij}}{N_{i}}$, where $X_{ij}$ represents the transcript count of gene $j$ in cell $i$, and $N_{j}$ is the total transcript count for cell $i$.
One can extend this concept to a Markov chain model (Teschendorff and Enver, 2017) to consider the interaction between the genes. Assume there is a predefined graph representing the gene-gene interaction (e.g., protein-protein interaction (PPI) network) from the existing database. The transition probability between genes $i$ and $j$ in cell $c$ is
$p_{ij}^{(c)}=\frac{x_{j}^{(c)}}{\sum_{k\in N(i)}x_{k}^{(c)}}=\frac{x_{j}^{(c)}}{(\mathbf{A}\mathbf{x}^{(c)})_{i}},$
where $x_{i}^{(c)}$ is the expression level of gene $i$ in cell $c$, $N(i)$ are the neighbors of gene $i$ in the graph, and $\mathbf{A}$ is the adjacency matrix of the graph. The corresponding stationary distribution is
$\pi^{(c)}_{i}=\frac{x^{(c)}_{i}(\mathbf{A}\mathbf{x}^{(c)})_{i}}{\mathbf{x}^{(c)^{T}}\mathbf{A}\mathbf{x}^{(c)}},$
@@
where $\widetilde{E}$ includes all edges on the graph. The entropy of cell $c$ is given by $\text{MCE}^{(c)}$, computed using $\pi_{i}^{(c)}$ and $p_{ij}^{(c)}$. To determine the weights, (Shi et al., 2020) proposes to optimize interaction weights based on cell expression $\boldsymbol{\pi}^{(c)}=\frac{\mathbf{x}^{(c)}}{\|\mathbf{x}^{(c)}\|_{\mathcal{L}_{1}}}$. For each cell, its MCE is maximized by solving
@@ 3.1.2 Discrete Dynamics Modeling
Diffusion Pseudotime
Previous pseudotime methods, such as those based on geodesic paths or simple assumptions of pseudotime, typically lacked an underlying dynamical model to explain cell state transitions. These methods often relied on the assumption of continuous trajectories without explicitly modeling the stochastic processes driving those transitions. In contrast, methods like DPT (Haghverdi et al., 2016) and PAGA (Wolf et al., 2019)introduce stochastic dynamics through Markov chains. While still estimating pseudotime, these methods incorporate random walk-defined observed samples of single cells, allowing for a more quantitative treatment of how cell states evolve over time, with transitions captured probabilistically. As a result, they provide a more mechanistic approach to pseudotime estimation, making them a natural progression from traditional geodesic-based methods.
Motivated by the DiffusionMap (Coifman and Lafon, 2006) algorithm for dimensionality reduction, DPT constructs a Markov chain between cells, defines a distance metric, and uses this distance as pseudotime. The transition probability of cell $i$ moving to cell $j$ is computed using a simple Gaussian kernel, defined as
$T_{ij}=\frac{1}{Z}W_{ij}=\frac{1}{Z}\left(\frac{2\sigma_{i}\sigma_{j}}{\sigma_{i}^{2}+\sigma_{j}^{2}}\right)\exp\left(-\frac{\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}}{2(\sigma_{i}^{2}+\sigma_{j}^{2})}\right),$
where $Z_{i}=\sum_{j\in\mathcal{N}(i)}W_{ij}$ is the normalization factor, and the hyperparameters $\sigma_{i},\sigma_{j}$ are the Gaussian kernel widths for cell $i$ and cell $j$. DPT assumes that the distance in the eigenspace of the transition matrix $T$ is related to the pseudotime ordering of cells. After removing steady-state eigenspace, the system’s dynamics are captured by the transition matrix $\bar{\mathbf{T}}=\mathbf{T}-\mathbf{\psi}_{1}\mathbf{\psi}_{1}^{T}$, where $\mathbf{\psi}_{1}$ represent the eigenvector corresponding to the largest eigenvalue of the transition matrix. The dynamics are analyzed by summing all $t$-step transition matrices to compute the cumulative probability of state transitions across multiple walk lengths
$\mathbf{M}=\sum_{t=1}^{\infty}\bar{\mathbf{T}}^{t}=(1-\bar{\mathbf{T}})^{-1}-\mathbf{I}.$
Using this matrix $\mathbf{M}$, a new distance metric is defined as
$\text{dpt}^{2}(i,j)=\|\mathbf{M}_{i,\cdot}-\mathbf{M}_{j,\cdot}\|=\sum_{k=2}^{n}\left(\frac{\lambda_{k}}{1-\lambda_{k}}\right)^{2}({\psi}_{k}(i)-{\psi}_{k}(j))^{2},$
@@
Random Walk with Directionality
Building on the inferred DPT, Palantir (Setty et al., 2019) introduces directionality into the cellular random walk, which is further developed in (Stassen et al., 2021; Pandey and Zafar, 2022). One simple approach is to prune the weight matrix as follows:
@@
For the directional Markov Chain induced by the weight matrix $\bar{\mathbf{W}}$, terminal states can be determined from its stationary distribution. Using absorption Markov Chain theory, a cell fate matrix $\mathbf{F}$ can be derived from canonical form of transition probability matrix in Equation (2.1), where $\mathbf{F}=(\mathbf{I}-\mathbf{Q})^{-1}\mathbf{S}$. Here, the element $\mathbf{F}_{ij}$ represents the probability that a random walk starting from transient cell $i$ will eventually be absorbed by terminal cell $j$. Specifically, the fate vector $\mathbf{f}_{i}$ of a transient cell $i$ corresponds to the $i$-th row of $\mathbf{F}$, capturing the probability of differentiation into various states. To quantify the differentiation potential of a cell, one can then the entropy of the fate vector, or the Kullback-Leibler (KL) divergence between the fate vector $\mathbf{f}_{i}$ and the average fate vector $\mathbf{\bar{f}}$.
Another method to define the directional random walk on cells is Population Balance Analysis (PBA) (Weinreb et al., 2018). Let $\mathbf{G}$ be the $k$-nearest neighbor graph of $\{\mathbf{X}\}$ and $\mathbf{L}$ its graph Laplacian. A potential function is defined by $\mathbf{V}=\dfrac{1}{2}\mathbf{L}_{N}^{\dagger}\mathbf{R}$ , where $\mathbf{L}_{N}^{\dagger}$ denotes the pseudo-inverse of $L_{N}$ and $\mathbf{R}$ is the estimated cell population production rate vector at each node, using the gene expression of predefined lists of proliferation relevant genes. The transition probabilities of this Markov chain are then directed by the potential function
$P_{ij}=\begin{cases}\exp\left(\dfrac{\mathbf{V}_{i}-\mathbf{V}_{j}}{D}\right)\quad&\text{if}\ (i,j)\ \text{is in }\mathbf{G}_{N}\\ 0&\text{otherwise}\end{cases}$
@@
Dissecting Dynamical Structure
While previous methods such as DPT, Palantir and PBA construct the random walk dynamics on snapshots of individual cells, methods like MuTrans (Zhou et al., 2021b) and CellRank (Lange et al., 2022; Weiler et al., 2024) take a deeper look into the dynamical structure of the system itself, especially focusing on metastability and attractor structures, therefore robustly dissecting the system’s latent dynamics to identify long-term patterns of the cell-state transition.
MuTrans (Zhou et al., 2021b) adopts a multi-scale reduction technique for a diffusion-based, unidirectional cellular random walk, to infer stable and transient cells from snapshot scRNA-seq. Central to MuTrans is a membership matrix $\chi_{i,k}$ representing the soft clustering probability that cell $i$ belongs to attractor $S_{k},k=1,..,K$, which could be interpreted as cell cluster. A transient cell might have multiple positive components in its attractor membership, while the distribution of stable cells tends to be concentrated in one specific attractor. Meanwhile, MuTrans also reduces the dynamics on the attractor level, using $\mathbf{P}^{(\text{coar})}\in\mathbb{R}^{K\times K}$ represent the coarsened transition matrix, and $\boldsymbol{\pi}^{(\text{coar})}$ represent the stationary distribution of the coarse-grained Markov chain. The original cell-to-cell dynamics can then be reconstructed from the coarsened dynamics, and the transition probability between cells $i$ and $j$ is given by
@@
The goal is to minimize the discrepancy between the reconstructed cell-cell dynamics $\hat{\mathbf{P}}$ and the actual dynamics $\mathbf{P}$, which is achieved by minimizing $\|\hat{\mathbf{P}}-\mathbf{P}\|^{2}$. This can be done using an EM-like algorithm, alternately optimizing the elements of $\mathbf{P}^{(\text{coar})}$ and $\chi_{i,j}$. With the inferred attractor membership matrix and coarse-grained transition probabilities among clusters, MuTrans then constructs a dynamical manifold inspired by the energy landscape concept (Zhou and Li, 2016) to visualize the transient and stable cells, and uses transition path theory (Vanden-Eijnden et al., 2010) based on ${P}^{(\text{coar})}$ to calculate the most probable transition paths among attractors.
CellRank extends the analysis by introducing a coarse-graining strategy for directed cellular random walks, such as induced by pseudotime (Setty et al., 2019) or RNA velocity described in Section 5.1. The approach begins with the clustering of cells into macro-states (i.e. attractors) using the GPCCA (Generalized Perron Cluster Cluster Analysis) algorithm (Reuter et al., 2019), which is based on the Shur decomposition of the directed transition matrix $\mathbf{P}$. The membership matrix $\mathbf{\chi}$ and the coarse-grained transition probability matrix between attractors $\mathbf{P}^{(\text{coar})}$ can be computed based on the decomposition. Once the attractors are identified, the terminal states can be determined in which the diagonal elements of the coarse transition matrix $\mathbf{P}^{(\text{coar})}$ exceed a certain threshold. Cells in terminal states can then be treated as absorption sets of the random walk, and the cell fate vector could be computed similarly to Palantir.

### v2

@@ 3 Dynamic Modeling of Single-Cell Transcriptomics
In this section, we describe the dynamical systems models for scRNA-seq datasets. We begin with methods for snapshot data, i.e., cells sequenced at a single time point, including pseudotime methods, discrete Markov chain methods, and continuous RNA velocity methods and their extensions. Next, we review methods targeted for temporally resolved scRNA-seq data, majorly based on various formulations and extensions of optimal transport (OT) based methods.
@@ 3.1 Snapshot Single-Cell RNA-Seq
When analyzing such a snapshot of “cell state ensembles”, several approaches have been developed to uncover the underlying dynamical processes. One popular type of method, pseudotime, ranks individual cells temporally based on the structure of the data manifold or prior biological knowledge. Other techniques focus on modeling stochastic dynamics over the point clouds of observed cells, yielding discrete random walk analyses. Additionally, continuous differential equation models have been proposed to infer the data-generating process of snapshot scRNA-seq dataset, to use these models for future dynamical predictions. In the following sections, we will explore these methods in more detail. An overview of these approaches is provided in Figure 2.
{adjustwidth}
[…]
@@ 3.1.1 Pseudotime Methods
Given snapshot data from single-cell sequencing, where the data matrix $\textbf{X}\in\mathbb{R}^{n\times d}$, pseudotime assigns a positive real number for each cell to reflect its order during a dynamical process. Let $\boldsymbol{x}_{i}\in\mathbb{R}^{d}$ represent the state vector (such as mRNA expression) for the $i$-th cell. The pseudotime $t_{i}\in\mathbb{R}$ is then a mapping from the state vector $\boldsymbol{x}_{i}$ to a real number, i.e., $\boldsymbol{x}_{i}\mapsto t_{i}$.
Pseudotime can be viewed from two perspectives. First, the geodesic perspective considers that the cell’s state is constrained by a limited number of biological pathways, limiting the evolution of the state to a low-dimensional manifold embedded in the high-dimensional gene expression space. Given an initial state $\boldsymbol{x}_{0}$, the task of finding the mapping $\boldsymbol{x}_{i}\mapsto t_{i}$ becomes equivalent to determining the length of the evolutionary path between $\boldsymbol{x}_{0}$ and $\boldsymbol{x}_{i}$ along this manifold. Several tools, such as Monocle, Slingshot, DPT, and PAGA, have been developed based on this idea. Second, the entropy perspective recognizes that during natural biological development, as cells differentiate, they become more specialized and lose their potential for further differentiation.
Geodesic-Based Pseudotime Geodesic -based pseudotime aims to reconstruct these trajectories by leveraging graph-based methods and principal curve algorithms Deconinck et al. (2021). Two widely used approaches include Monocle Qiu et al. (2017); Cao et al. (2019) and Slingshot Street et al. (2018), along with numerous other methods. Monocle focuses on ordering cells using a minimum spanning tree and a refined PQ tree approach, and Slingshot can handle multiple lineages and smooth pseudotime across branching events. Both methods offer insights into cellular dynamics, helping to uncover the paths cells take through different states.
Monocle estimates pseudotime through two main steps: (1) ordering cells and (2) assigning pseudotime values. First, each cell’s state is reduced to a $d$-dimensional vector $\boldsymbol{x}_{i}\in\mathbb{R}^{d}$ using Independent Component Analysis (ICA). A complete graph is created where vertices represent cells, and edges are weighted by Euclidean distance. Cell ordering relies on the minimum spanning tree (MST) of this graph, which is refined using a PQ tree to mitigate noise from sequencing. The tree is constructed by first identifying the longest path (diameter path), classifying vertices as decisive or indecisive, and recursively building the tree by ordering decisive vertices and handling indecisive vertices through new P nodes. Once ordered, pseudotime is calculated as
$t(\boldsymbol{x}_{i})=t(\boldsymbol{x}_{\text{Parent}(i)})+\|\boldsymbol{x}_{i}-\boldsymbol{x}_{\text{Parent}(i)}\|,$
@@
Slingshot estimates pseudotime across multiple lineages. It begins by clustering cells into $K$ clusters and identifying lineages using the MST. After constructing the MST, a new lineage is formed at each branching point. Pseudotime is assigned using the principal curve algorithm, which involves projecting cells onto the curve, computing arc lengths, and smoothing iteratively. To handle inconsistent pseudotime across multiple lineages, Slingshot modifies the standard principal curve approach. It initializes the curve for each lineage through the centroids of its clusters, assigns weights to cells in multiple lineages based on projection distances, and constructs an average curve for smooth transitions at shared cell regions. The average curve is defined as
$\boldsymbol{c}_{\text{avg}}(t)=\frac{1}{M}\sum_{m=1}^{M}\boldsymbol{c}_{m}(t),$
where $\boldsymbol{c}_{m}$ is the principal curve of the $m$-th lineage at the branching point. The shrinkage process is defined as
$\boldsymbol{c}_{m}^{\text{new}}(t)=w_{m}(t)\boldsymbol{c}_{\text{avg}}(t)+(1-w_{m}(t))\boldsymbol{c}_{m}(t),$
@@
Entropy-Based Pseudotime One major challenge of the geodesic-based pseudotime is the appropriate determination of root cells, which often relies on prior biological knowledge. From a physical understanding, a cell’s pseudotime reflects the directionality of the underlying dynamical process, which the concept of entropy could quantify. Heuristically, higher entropy values typically indicate a more undifferentiated or pluripotent state where genes are more randomly expressed and the association between genes could be more prevalent. In comparison, lower entropy values suggest differentiated states, where the gene expression profile could be more concentrated on only a small number of pathways, and the gene interaction network could be more modular Gandrillon et al. (2021). A proper entropy score based on such intuitions can thus be leveraged to estimate a cell’s relative position along developmental trajectories. Several methods have been developed from this perspective Grün et al. (2016); Guo et al. (2017); Teschendorff and Enver (2017); Shi et al. (2020); Jin et al. (2018); Liu et al. (2020).
As a simple implementation, the entropy for one cell $i$ is defined as Grün et al. (2016)
@@
where $p_{ij}=\frac{X_{ij}}{N_{i}}$, and $X_{ij}$ represents the transcript count of gene $j$ in cell $i$, and $N_{j}$ is the total transcript count for cell $i$.
One can extend this concept to a Markov chain model Teschendorff and Enver (2017) to consider the interaction between the genes. Assume there is a predefined graph representing the gene–gene interaction (e.g., protein–protein interaction (PPI) network) from the existing database. The transition probability between genes $i$ and $j$ in cell $c$ is
$p_{ij}^{(c)}=\frac{x_{j}^{(c)}}{\sum_{k\in N(i)}x_{k}^{(c)}}=\frac{x_{j}^{(c)}}{(\boldsymbol{A}\boldsymbol{x}^{(c)})_{i}},$
where $x_{i}^{(c)}$ is the expression level of gene $i$ in cell $c$, $N(i)$ are the neighbors of gene $i$ in the graph, and $\boldsymbol{A}$ is the adjacency matrix of the graph. The corresponding stationary distribution is
$\pi^{(c)}_{i}=\frac{x^{(c)}_{i}(\boldsymbol{A}\boldsymbol{x}^{(c)})_{i}}{\boldsymbol{x}^{(c)^{T}}\boldsymbol{A}\boldsymbol{x}^{(c)}},$
@@
where $\widetilde{E}$ includes all edges on the graph. The entropy of cell $c$ is given by $\text{MCE}^{(c)}$, computed using $\pi_{i}^{(c)}$ and $p_{ij}^{(c)}$. To determine the weights, ref. Shi et al. (2020) proposes to optimize interaction weights based on cell expression $\boldsymbol{\pi}^{(c)}=\frac{\boldsymbol{x}^{(c)}}{\|\boldsymbol{x}^{(c)}\|_{\mathcal{L}_{1}}}$. For each cell, its MCE is maximized by solving
@@ 3.1.2 Discrete Dynamics Modeling
Diffusion Pseudotime Previous pseudotime methods, such as those based on geodesic paths or simple assumptions of pseudotime, typically lacked an underlying dynamical model to explain cell state transitions. These methods often relied on the assumption of continuous trajectories without explicitly modeling the stochastic processes driving those transitions. In contrast, methods like DPT Haghverdi et al. (2016) and PAGA Wolf et al. (2019) introduce stochastic dynamics through Markov chains. While still estimating pseudotime, these methods incorporate random walk-defined observed samples of single cells, allowing for a more quantitative treatment of how cell states evolve over time, with transitions captured probabilistically. As a result, they provide a more mechanistic approach to pseudotime estimation, making them a natural progression from traditional geodesic-based methods.
Motivated by the DiffusionMap Coifman and Lafon (2006) algorithm for dimensionality reduction, DPT constructs a Markov chain between cells, defines a distance metric, and uses this distance as pseudotime. The transition probability of cell $i$ moving to cell $j$ is computed using a simple Gaussian kernel, defined as
$T_{ij}=\frac{1}{Z}W_{ij}=\frac{1}{Z}\left(\frac{2\sigma_{i}\sigma_{j}}{\sigma_{i}^{2}+\sigma_{j}^{2}}\right)\exp\left(-\frac{\|\boldsymbol{x}_{i}-\boldsymbol{x}_{j}\|^{2}}{2(\sigma_{i}^{2}+\sigma_{j}^{2})}\right),$
where $Z_{i}=\sum_{j\in\mathcal{N}(i)}W_{ij}$ is the normalization factor, and the hyperparameters $\sigma_{i},\sigma_{j}$ are the Gaussian kernel widths for cell $i$ and cell $j$. DPT assumes that the distance in the eigenspace of the transition matrix $T$ is related to the pseudotime ordering of cells. After removing steady-state eigenspace, the system’s dynamics are captured by the transition matrix $\bar{\boldsymbol{T}}=\boldsymbol{T}-\boldsymbol{\psi}_{1}\boldsymbol{\psi}_{1}^{T}$, where $\boldsymbol{\psi}_{1}$ represent the eigenvector corresponding to the largest eigenvalue of the transition matrix. The dynamics are analyzed by summing all $t$-step transition matrices to compute the cumulative probability of state transitions across multiple walk lengths
$\boldsymbol{M}=\sum_{t=1}^{\infty}\bar{\boldsymbol{T}}^{t}=(1-\bar{\boldsymbol{T}})^{-1}-\boldsymbol{I}.$
Using this matrix $\boldsymbol{M}$, a new distance metric is defined as
$\text{dpt}^{2}(i,j)=\|\boldsymbol{M}_{i,\cdot}-\boldsymbol{M}_{j,\cdot}\|=\sum_{k=2}^{n}\left(\frac{\lambda_{k}}{1-\lambda_{k}}\right)^{2}({\psi}_{k}(i)-{\psi}_{k}(j))^{2},$
@@
Random Walk with Directionality Building on the inferred DPT, Palantir Setty et al. (2019) introduces directionality into the cellular random walk, which is further developed in Stassen et al. (2021); Pandey and Zafar (2022). One simple approach is to prune the weight matrix as follows:
@@
For the directional Markov chain induced by the weight matrix $\bar{\boldsymbol{W}}$, terminal states can be determined from its stationary distribution. Using absorption Markov chain theory, a cell fate matrix $\boldsymbol{F}$ can be derived from the canonical form of transition probability matrix in Equation (1), where $\boldsymbol{F}=(\boldsymbol{I}-\boldsymbol{Q})^{-1}\boldsymbol{S}$. Here, the element $\boldsymbol{F}_{ij}$ represents the probability that a random walk starting from transient cell $i$ will eventually be absorbed by terminal cell $j$. Specifically, the fate vector $\boldsymbol{f}_{i}$ of a transient cell $i$ corresponds to the $i$-th row of $\boldsymbol{F}$, capturing the probability of differentiation into various states. To quantify the differentiation potential of a cell, one can then determine the entropy of the fate vector, or the Kullback–Leibler (KL) divergence between the fate vector $\boldsymbol{f}_{i}$ and the average fate vector $\boldsymbol{\bar{f}}$. Author: Please check that the intended meaning has been retained.
Another method to define the directional random walk on cells is Population Balance Analysis (PBA) Weinreb et al. (2018). Let $\boldsymbol{G}$ be the $k$-nearest neighbor graph of $\{\boldsymbol{X}\}$ and $\boldsymbol{L}$ its graph Laplacian. A potential function is defined by $\boldsymbol{V}=\dfrac{1}{2}\boldsymbol{L}_{N}^{\dagger}\boldsymbol{R}$, where $\boldsymbol{L}_{N}^{\dagger}$ denotes the pseudo-inverse of $L_{N}$ and $\boldsymbol{R}$ is the estimated cell population production rate vector at each node, using the gene expression of predefined lists of proliferation relevant genes. The transition probabilities of this Markov chain are then directed by the potential function
$P_{ij}=\begin{cases}\exp\left(\dfrac{\boldsymbol{V}_{i}-\boldsymbol{V}_{j}}{D}\right)\quad&\text{if}\ (i,j)\ \text{is in }\boldsymbol{G}_{N}\\ 0&\text{otherwise}\end{cases}$
@@
Dissecting Dynamical Structure While previous methods such as DPT, Palantir, and PBA construct the random walk dynamics on snapshots of individual cells, methods like MuTrans Zhou et al. (2021) and CellRank Lange et al. (2022); Weiler et al. (2024) take a deeper look into the dynamical structure of the system itself, especially focusing on metastability and attractor structures, therefore robustly dissecting the system’s latent dynamics to identify long-term patterns of the cell-state transition.
MuTrans Zhou et al. (2021) adopts a multi-scale reduction technique for a diffusion-based, unidirectional cellular random walk, to infer stable and transient cells from snapshot scRNA-seq. Central to MuTrans is a membership matrix $\chi_{i,k}$ representing the soft clustering probability that cell $i$ belongs to attractor $S_{k},k=1,\ldots,K$, which could be interpreted as a cell cluster. A transient cell might have multiple positive components in its attractor membership, while the distribution of stable cells tends to be concentrated in one specific attractor. Meanwhile, MuTrans also reduces the dynamics on the attractor level, using $\boldsymbol{P}^{(\text{coar})}\in\mathbb{R}^{K\times K}$ to represent the coarsened transition matrix, and $\boldsymbol{\pi}^{(\text{coar})}$ to represent the stationary distribution of the coarse-grained Markov chain. The original cell-to-cell dynamics can then be reconstructed from the coarsened dynamics, and the transition probability between cells $i$ and $j$ is given by
@@
The goal is to minimize the discrepancy between the reconstructed cell–cell dynamics $\hat{\boldsymbol{P}}$ and the actual dynamics $\boldsymbol{P}$, which is achieved by minimizing $\|\hat{\boldsymbol{P}}-\boldsymbol{P}\|^{2}$. This can be done using an EM-like algorithm, alternately optimizing the elements of $\boldsymbol{P}^{(\text{coar})}$ and $\chi_{i,j}$. With the inferred attractor membership matrix and coarse-grained transition probabilities among clusters, MuTrans then constructs a dynamical manifold inspired by the energy landscape concept Zhou and Li (2016) to visualize the transient and stable cells, and uses transition path theory Vanden-Eijnden et al. (2010) based on ${P}^{(\text{coar})}$ to calculate the most probable transition paths among attractors.
CellRank extends the analysis by introducing a coarse-graining strategy for directed cellular random walks, such as those induced by pseudotime Setty et al. (2019) or RNA velocity as described in Section 5.1. The approach begins with the clustering of cells into macro-states (i.e., attractors) using the GPCCA (Generalized Perron Cluster Cluster Analysis) algorithm Reuter et al. (2019), which is based on the Shur decomposition of the directed transition matrix $\boldsymbol{P}$. The membership matrix $\boldsymbol{\chi}$ and the coarse-grained transition probability matrix between attractors $\boldsymbol{P}^{(\text{coar})}$ can be computed based on the decomposition. Once the attractors are identified, the terminal states can be determined in which the diagonal elements of the coarse transition matrix $\boldsymbol{P}^{(\text{coar})}$ exceed a certain threshold. Cells in terminal states can then be treated as absorption sets of the random walk, and the cell fate vector could be computed similarly to Palantir.